# XLAVS-R: Cross-Lingual Audio-Visual Speech Representation Learning for Noise-Robust Speech Perception

### v1

@@ Abstract
[…]
@@ A AV A AV
AVFormer Seo et al. (2023) 60K 0 ${}^{\star}$ 0
FAVA May et al. (2023) 2.8K 0 ${}^{\dagger}$ 0
@@
AV-HuBERT Shi et al. (2022a) 0 1.8K 0 ${}^{\star}$
AV-data2vec Lian et al. (2024) 0 1.8K 0 ${}^{\star}$
@@
u-HuBERT Hsu and Shi (2022) 0.5K 1.8K ${}^{\star}$ ${}^{\star}$
VATLM Zhu et al. (2024) 4.3K 1.8K ${}^{\star}$ ${}^{\star}$
@@
Table 1: Pre-training data type, amount and language coverage in audio-visual speech perception models (A: audio, AV: audio-visual). Compared to prior work, XLAVS-R exploits audio-only speech data for efficient data scaling and language coverage expansion. ${}^{\star}$ English-only.
@@
We conduct extensive evaluations on the MuAViC benchmark (Anwar et al., 2023). XLAVS-R yields SOTA performance on speech recognition in 9 languages and speech-to-text translation in 6 X-to-English pairs (§5.1, §5.2), notably improving their robustness to noisy inputs. Furthermore, XLAVS-R enables zero-shot audio-visual ability in downstream tasks with audio-only fine-tuning (FT): fine-tuning our XLAVS-R 2B model without supervised AV downstream data achieves the best audio-visual performances in noisy settings even compared to the corresponding audio-visual supervised fine-tuned model (§5.4). This confirms the benefits of our pre-training strategy and relaxes the data requirements for downstream tasks, making it possible to tackle AVSR and AVSTT tasks without labeled—transcribed or translated—audio-visual data.
@@ 3.4 Single-Round AV Training with Self-Supervised Audio-Only Targets
Clean environment, Test WER (%) normal-$\downarrow$
@@
Noise environment, Test WER (%) normal-$\downarrow$
@@ 4 Experiments
Data.
For AV pre-training, we leverage a total 1.2K hours of data in 9 languages: English (En), Arabic (Ar), German (De), Greek (El), Spanish (Es), French (Fr), Italian (It), Portuguese (Pt) and Russian (Ru). For A-only pre-training, we follow Babu et al. (2022) for training on 436K-hour data in 128 languages.
@@
Clean environment, Test BLEU normal-$\uparrow$
@@
Noisy environment, Test BLEU normal-$\uparrow$
@@
We build XLS-R (§3.1) and XLAVS-R (§3) models in two model sizes: 300M and 2B. The number of encoder layers/embedding dimension/feed forward dimension are 24/1024/4096 and 48/1920/7680 respectively. For audio-visual training targets of XLAVS-R, we extract audio-only speech representation from the 36th layer of XLS-R 2B and quantize it with 2000 k-means clusters. For fine-tuning AVSR and AVS2TT models, we follow Anwar et al. (2023) to add a Transformer decoder that has 6 layers, an embedding dimension of 768, and feed forward network dimension of 3072. The visual feature extractor of our model is modified ResNet-18 as in AV-HuBERT and prior lipreading works (Stafylakis and Tzimiropoulos, 2017; Martinez et al., 2020; Ma et al., 2021). For XLAVS-R, we set both $p_{m}$ and $p_{a}$ to 0.5 for pre-training which is the same with AV/u-HuBERT. For fine-tuning all pre-trained models, we set $p_{m}$ and $p_{a}$ to 0.5 and 0, where the models use 50% AV modality and 50% A-only modality.
Noise Augmentation.
Following Shi et al. (2022b) and Anwar et al. (2023), we randomly augment the input samples with additive noises at 0dB of signal-to-noise ratio. Noise audio clips are sampled from the MUSAN dataset Snyder et al. (2015) as well as drawn from MuAViC train set (for babble type of noises). We augment 25% of the input in pre-training and 50% in fine-tuning with noises.
Evaluation.
We evaluate AVSR and AVS2TT in audio-only (“A”) and audio-visual (“AV”) test modes, where the former leverages only audio modality in inference while the latter leverages both audio and visual modalities. We select the best checkpoint by validation accuracy for inference. We use beam search with default hyper-parameters and beam size of 5. For AVSR, we apply Whisper text normalizer Radford et al. (2023) before calculating WER (word error rate). For AVS2TT, we compute BLEU  Papineni et al. (2002) using SacreBLEU Post (2018) with default options and the built-in 13a tokenizer. For the simulation of noisy test environment, we add babble type of noises drawn from MuAViC test set.
@@ 5.1 Multilingual Speech Recognition
Overall, XLAVS-R models outperform the two baselines by a large margin on average in clean and noisy settings, while remaining competitive with English-only pre-trained models in English. Results of experiments with additional AV data for domain and language coverage during AV pre-training are available in Appendix A and Table 5.
@@ 5.2 Multilingual Speech-To-Text Translation
Overall, these results show that our XLAVS-R models outperform two baselines by a large margin in all translation directions for both clean and noisy settings, both A-only and AV test modes, and both in-domain and OOD. Results of experiments with additional AV data for domain and language coverage during AV pre-training are available in Appendix A and Table 6.
@@ 5.3 Ablation Experiments of XLAVS-R
In Figure 2111Appendix Table 7 contains the exact numbers plotted., we validate the effectiveness of the key changes from AV-HuBERT to XLAVS-R, step-by-step starting from AV-HuBERT pre-trained on MuAViC-En. First, simplifying the training process – single-round AV-HuBERT training with self-supervised contextualized audio-only units from XLS-R (+ Single-round w/SSL units) instead of the original 5 training rounds – actually yields moderate improvements. Second, switching to a learnable audio feature extractor (+ Learnable AFE) shows the biggest improvements in both clean and noisy settings, especially in the low-resource languages (El and Ru). Third, augmenting AV pre-training with multilingual data (+ MuAViC non-English) is the second-highest improvement indicating that multilingual audio-visual representation learning is critical in the self-supervised pre-training stage. Last, the introduction of the audio-only pre-training stage and multilingual audio-only resources (+ A-only pre-training (XLAVS-R)) further improves both A and AV performance in the in-domain evaluations. It yields even higher improvements in OOD evaluations, suggesting that audio-only pretraining may enhance robustness against domain shift.
@@ 5.4 Zero-shot Audio-Visual Inference
6 Conclusion
We introduced XLAVS-R, a cross-lingual audio-visual speech representation for noise-robust speech perception in over 100 languages. Since audio-visual data is harder to come by than audio-only data, we injected the visual modality by continuing training an audio-only pre-trained model. We also design a simpler yet more effective training scheme and improved architecture compared to previous state-of-the-art models. Extensive evaluation on the MuAViC benchmark shows XLAVS-R achieves SOTA performance on both audio-visual speech recognition and translation tasks, yielding particularly large improvements in noisy settings. We also show that XLAVS-R effectively leverages unlabeled audio-only multilingual speech data and audio-visual data in self-supervised learning to show enhanced zero-shot audio-visual ability for downstream tasks.
Limitations
Appendix A Results with Additional Training Data
We also experiment with additional data for domain and language coverage during the audio-visual self-supervised learning phase. On top of the existing 1.2K-hour data, we add 7.1K hours of in-house AV data in 100+ languages and train XLAVS-R with a total of 8.3K hours in 100+ languages, as summarized in the last row of Table 4. We show the AVSR and AVS2TT results comparing the XLAVS-R 300M models trained on different sizes and numbers of the language of the training dataset.
Table 6 shows AVS2TT performances of 6 X-to-English pairs on OOD and in-domain test sets, where the trend of the prominent improvement in OOD is similar to that of AVSR comparison. Additional AV SSL dataset improves the average performance of in-domain noisy setting by 0.7 BLEU for A-only test mode and 0.4 BLEU for AV test mode, while slightly degrades in clean settings by 0.1 BLEU in A-only test mode and 0.2 in AV test mode on average. In out-of-domain evaluation, augmented data in AV SSL enhances the translation quality by up to 1.5 BLEU on average in noisy environments.
Hours Languages
A AV A AV
Audio-Only Pre-Training Data
AVFormer Seo et al. (2023) 60K 0 ${}^{\star}$ 0
FAVA May et al. (2023) 2.8K 0 ${}^{\dagger}$ 0
Audio-Visual Pre-Training Data
AV-HuBERT Shi et al. (2022a) 0 1.8K 0 ${}^{\star}$
AV-data2vec Lian et al. (2024) 0 1.8K 0 ${}^{\star}$
AV2AV Choi et al. (2023) 0 7K 0 ${}^{\dagger}$
Audio-Only & Audio-Visual Pre-Training Data
u-HuBERT Hsu and Shi (2022) 0.5K 1.8K ${}^{\star}$ ${}^{\star}$
VATLM Zhu et al. (2024) 4.3K 1.8K ${}^{\star}$ ${}^{\star}$
XLAVS-R (w/ basic AV data) 436K 1.2K 128 9
XLAVS-R (w/ extended AV data) 436K 8.3K 128 ${}^{\dagger}$
Table 4: Pre-training data type, amount and language coverage in audio-visual speech perception models (A: audio, AV: audio-visual). Compared to prior work, XLAVS-R exploits audio-only speech data for efficient data scaling and language coverage expansion. ${}^{\star}$ English-only. ${}^{\dagger}$ Estimated by a speech language identification model.
@@ A AV A AV
Clean environment, Test WER (%) normal-$\downarrow$
@@
Noise environment, Test WER (%) normal-$\downarrow$
@@
Table 5: Comparison of XLAVS-R 300M models trained on different sizes of the audio-visual pre-training dataset. Results show the performance of audio-visual speech recognition (AVSR) on in-domain (MuAViC) and out-of-domain (OOD, FLEURS) evaluation in test mode of audio-only (A) and audio-video (AV).
@@
Clean environment, Test BLEU normal-$\uparrow$
@@
Noisy environment, Test BLEU normal-$\uparrow$
@@
Table 6: Comparison of XLAVS-R 300M models trained on different sizes of the audio-visual pre-training dataset. Results show the performance of audio-visual speech-to-text translation (AVS2TT) on in-domain (MuAViC) and out-of-domain (OOD, FLEURS) evaluation in test mode of audio-only (A) and audio-video (AV).
@@
Clean environment, Test WER (%) normal-$\downarrow$
@@
Noisy environment, Test WER (%) normal-$\downarrow$
@@
Clean environment, Test WER (%) normal-$\downarrow$
@@
Noisy environment, Test WER (%) normal-$\downarrow$

### v2

@@ Abstract
HyoJung HanAthanks:  Work done at Meta AI Mohamed Anwar11footnotemark: 1 Juan PinoV Wei-Ning HsuV Marine CarpuatABowen ShiVChanghan WangV AUniversity of Maryland, USA VMeta AI, USA hjhan@cs.umd.educhanghan@meta.com
@@ A AV A AV
AVFormer Seo et al. (2023) 60K 0 1 0
FAVA May et al. (2023) 2.8K 0 $>$100 0
@@
AV-HuBERT Shi et al. (2022a) 0 1.8K 0 1
AV-data2vec Lian et al. (2024) 0 1.8K 0 1
@@
u-HuBERT Hsu and Shi (2022) 0.5K 1.8K 1 1
VATLM Zhu et al. (2024) 4.3K 1.8K 1 1
@@
XLAVS-R (+ extended AV data) 436K 8.3K 128 100+
Table 1: Pre-training data type, amount and language coverage in audio-visual speech perception models (A: audio, AV: audio-visual). Compared to prior work, XLAVS-R exploits audio-only speech data for efficient data scaling and language coverage expansion. English-only.
@@
We conduct extensive evaluations on the MuAViC benchmark (Anwar et al., 2023). XLAVS-R yields SOTA performance on speech recognition in 9 languages and speech-to-text translation in 6 X-to-English pairs (§5.1, §5.2), notably improving their robustness to noisy inputs. Furthermore, XLAVS-R enables zero-shot audio-visual ability in downstream tasks with audio-only fine-tuning (FT): fine-tuning our XLAVS-R 2B model without supervised AV downstream data achieves the best audio-visual performances in noisy settings even compared to the corresponding audio-visual supervised fine-tuned model (§5.4). This confirms the benefits of our pre-training strategy and relaxes the data requirements for downstream tasks, making it possible to tackle AVSR and AVS2TT tasks without labeled—transcribed or translated—AV data.
@@ 3.4 Single-Round AV Training with Self-Supervised Audio-Only Targets
Clean environment, Test WER (%) $\downarrow$
@@
Noise environment, Test WER (%) $\downarrow$
@@ 4 Experiments
Data. For AV pre-training, we leverage a total 1.2K hours of data in 9 languages: English (En), Arabic (Ar), German (De), Greek (El), Spanish (Es), French (Fr), Italian (It), Portuguese (Pt) and Russian (Ru). We also experiment with additional data for domain and language coverage. On top of the above 1.2K hour data, we add 7.1K hours of in-house AV data in 100+ languages and train XLAVS-R with a total of 8.3K hours in 100+ languages, as summarized in the last row of Table 1. The number of hours we use in the training of each experiment setup for 9 language we focus on are shown in Table 6.
For A-only pre-training, we follow Babu et al. (2022) for training on 436K-hour data in 128 languages.
@@
Clean environment, Test BLEU $\uparrow$
@@
Noisy environment, Test BLEU $\uparrow$
@@
We build XLS-R (§3.1) and XLAVS-R (§3) models in two model sizes: 300M and 2B. The number of encoder layers/embedding dimension/feed forward dimension are 24/1024/4096 and 48/1920/7680 respectively. For audio-visual training targets of XLAVS-R, we extract audio-only speech representation from the 36th layer of XLS-R 2B and quantize it with 2000 k-means clusters. For fine-tuning AVSR and AVS2TT models, we follow Anwar et al. (2023) to add a Transformer decoder that has 6 layers, an embedding dimension of 768, and feed forward network dimension of 3072. The visual feature extractor of our model is modified ResNet-18 as in AV-HuBERT and prior lipreading works (Stafylakis and Tzimiropoulos, 2017; Martinez et al., 2020; Ma et al., 2021). For XLAVS-R, we set both $p_{m}$ and $p_{a}$ to 0.5 for pre-training which is the same with AV/u-HuBERT. For fine-tuning all pre-trained models, we set $p_{m}$ and $p_{a}$ to 0.5 and 0, where the models use 50% AV modality and 50% A-only modality. We set $\alpha$ as 1.0 following the AV-HuBERT models [^1] https://facebookresearch.github.io/av_hubert.
Noise Augmentation. Following Shi et al. (2022b) and Anwar et al. (2023), we randomly augment the input samples with additive noises at 0dB of signal-to-noise ratio. Noise audio clips are sampled from the MUSAN dataset Snyder et al. (2015) as well as drawn from MuAViC train set for babble type. We augment 25% of the input in pre-training and 50% in fine-tuning with noises.
Evaluation. We evaluate AVSR and AVS2TT in audio-only (“A”) and audio-visual (“AV”) test modes, where the former leverages only audio modality in inference while the latter leverages both audio and visual modalities. We select the best checkpoint by validation accuracy for inference. We use beam search with default hyper-parameters and beam size of 5. For AVSR, we apply Whisper text normalizer Radford et al. (2023) before calculating WER (word error rate). For AVS2TT, we compute BLEU  Papineni et al. (2002) using SacreBLEU Post (2018) with default options and the built-in 13a tokenizer. For the simulation of noisy test environment, we add babble type of noises drawn from MuAViC test set.
@@ 5.1 Multilingual Speech Recognition
Overall, XLAVS-R models outperform the two baselines by a large margin on average in clean and noisy settings, while remaining competitive with English-only pre-trained models in English. Results of experiments with additional AV data for domain and language coverage during AV pre-training are available in Appendix 5.5 and Table 4. We also show that the findings are consistent with different types of noise in Appendix A, Table 7-8.
@@ 5.2 Multilingual Speech-To-Text Translation
Overall, these results show that our XLAVS-R models outperform two baselines by a large margin in all translation directions for both clean and noisy settings, both A-only and AV test modes, and both in-domain and OOD. Results of experiments with additional AV data for domain and language coverage during AV pre-training are available in Appendix 5.5 and Table 5.
@@ 5.3 Ablation Experiments of XLAVS-R
In Figure $2$Appendix Table 9 contains the exact numbers plotted., we validate the effectiveness of the key changes from AV-HuBERT to XLAVS-R, step-by-step starting from AV-HuBERT pre-trained on MuAViC-En. First, simplifying the training process – single-round AV-HuBERT training with self-supervised contextualized audio-only units from XLS-R (+ Single-round w/SSL units) instead of the original 5 training rounds – actually yields moderate improvements. Second, switching to a learnable audio feature extractor (+ Learnable AFE) shows the biggest improvements in both clean and noisy settings, especially in the low-resource languages (El and Ru). Third, augmenting AV pre-training with multilingual data (+ MuAViC non-English) is the second-highest improvement indicating that multilingual audio-visual representation learning is critical in the self-supervised pre-training stage. Last, the introduction of the audio-only pre-training stage and multilingual audio-only resources (+ A-only pre-training (XLAVS-R)) further improves both A and AV performance in the in-domain evaluations. It yields even higher improvements in OOD evaluations, suggesting that audio-only pretraining may enhance robustness against domain shift.
@@ 5.4 Zero-shot Audio-Visual Inference
Clean environment, Test WER (%) $\downarrow$
@@
Noise environment, Test WER (%) $\downarrow$
@@
Table 4: AVSR comparison of XLAVS-R 300M models trained on different sizes of the AV pre-training dataset.
@@
Clean environment, Test BLEU $\uparrow$
@@
Noisy environment, Test BLEU $\uparrow$
@@
Table 5: AVS2TT comparison of XLAVS-R 300M models trained on different sizes of the AV pre-training dataset.
5.5 Results with Additional Training Data
We also experiment with additional data for domain and language coverage during the audio-visual self-supervised learning phase.
Table 5 shows AVS2TT performances of six X-to-En pairs on OOD and in-domain test sets, where the trend of the prominent improvement in OOD is similar to that of AVSR comparison. Additional AV SSL dataset improves the average performance of in-domain noisy setting by 0.7 BLEU for A-only test mode and 0.4 BLEU for AV test mode, while slightly degrades in clean settings by 0.1 BLEU in A-only test mode and 0.2 in AV test mode on average. In out-of-domain evaluation, augmented data in AV SSL enhances the translation quality by up to 1.5 BLEU on average in noisy environments.
6 Conclusion
We introduced XLAVS-R, a cross-lingual audio-visual speech representation for noise-robust speech perception in over 100 languages. Since audio-visual data is harder to come by than audio-only data, we injected the visual modality by continuing training an audio-only pre-trained model. We also design a simpler yet more effective training scheme and improved architecture compared to previous state-of-the-art models. Extensive evaluation on the MuAViC benchmark shows XLAVS-R achieves SOTA performance on both audio-visual speech recognition and translation tasks, yielding particularly large improvements in noisy settings. We also show that XLAVS-R effectively leverages unlabeled audio-only multilingual speech data and audio-visual data in self-supervised learning resulting in enhanced zero-shot audio-visual ability for downstream tasks, and extended AV pre-training data augments robustness against domain shift.
Limitations
Appendix A Generalization to Different Types of Noise
As we mentioned in the Limitation section, the evaluation phase in the main tables is conducted on a single type of synthetic noise, Babble. To confirm the general effectiveness of XLAVS-R on noise robustness, we further experiment AVSR with different types of noise, Music and Noise from MUSAN dataset following Shi et al. (2022b).
Table 7 and 8 show the results in average and language-wise WER of each models on different types of noise. Overall, the noise robustness of Audio-Visual speech processing still holds with different kinds of noises. All AV models manage to get a higher quality of speech recognition with visual information than A-only mode in all new types of noise. Babble turns out to be the toughest noisy environment as shown in the above table, while Music and Noise are even able to achieve scores that are close to the clean environment with AV mode, which is consistent with Shi et al. (2022b).
Noise Type Clean Babble Music Noise
Mode A AV A AV A AV A AV
AV-HuBERT 35.7 35.0 73.5 56.1 45.8 41.3 45.8 41.3
u-HuBERT 35.6 35.0 71.0 55.8 45.1 41.0 45.4 41.0
XLAVS-R 300M 26.7 26.3 53.4 41.2 33.0 30.6 33.1 30.5
XLAVS-R 300M (extended) 26.5 26.1 53.9 41.0 32.7 30.1 33.0 30.3
XLAVS-R 2B 24.2 23.6 50.8 37.3 30.4 27.8 30.5 27.7
Table 7: Average WER of AVSR with different types of noise of Music and Noise from MUSAN dataset. The noise robustness of XLAVS-R is still the most effective with different kinds of noises.
Model Noise Type En Ar De El Es Fr It Pt Ru Avg
AV-HuBERT Noisy-A (Music) 6.7 97.3 65.0 57.8 30.1 29.7 34.4 35.8 55.8 45.8
Noisy-AV (Music) 2.9 94.1 59.8 53.4 24.5 26.1 28.8 29.8 51.9 41.3
Noisy-A (Noise) 7.3 96.1 65.5 58.4 29.2 29.4 34.0 36.3 56.3 45.8
Noisy-AV (Noise) 3.2 94.3 59.7 53.6 24.2 26.5 28.1 30.0 52.4 41.3
u-HuBERT Noisy-A (Music) 6.3 96.9 63.5 57.2 29.2 29.2 33.5 34.6 55.1 45.1
Noisy-AV (Music) 3.2 94.3 58.7 53.1 24.6 26.2 28.3 28.9 51.7 41.0
Noisy-A (Noise) 6.1 96.7 64.5 58.0 28.2 30.0 33.8 35.3 56.2 45.4
Noisy-AV (Noise) 2.9 93.5 59.1 53.9 23.8 26.6 28.1 29.2 52.0 41.0
XLAVS-R 300M Noisy-A (Music) 5.3 87.7 53.3 33.6 17.1 19.3 19.6 21.7 39.5 33.0
Noisy-AV (Music) 3.2 85.8 50.3 30.6 15.2 17.7 16.8 18.3 37.4 30.6
Noisy-A (Noise) 6.0 87.1 54.0 33.2 17.2 19.3 19.3 21.8 40.3 33.1
Noisy-AV (Noise) 3.5 84.8 50.7 30.0 14.8 17.9 16.4 18.1 38.0 30.5
XLAVS-R 300M (extended) Noisy-A (Music) 4.9 85.0 45.5 37.4 18.4 20.4 21.7 22.9 38.1 32.7
Noisy-AV (Music) 2.9 83.2 42.4 34.1 16.0 18.8 18.6 19.3 36.0 30.1
Noisy-A (Noise) 5.5 85.2 46.2 37.4 18.4 20.9 21.2 23.9 38.5 33.0
Noisy-AV (Noise) 2.9 83.1 42.8 34.2 15.7 19.3 17.9 20.0 36.5 30.3
XLAVS-R 2B Noisy-A (Music) 4.9 86.5 52.5 28.1 15.7 16.3 17.0 19.0 33.9 30.4
Noisy-AV (Music) 2.4 83.8 50.1 24.8 13.3 15.2 14.2 15.6 31.0 27.8
Noisy-A (Noise) 5.4 86.2 52.6 27.5 15.2 16.9 17.0 19.2 34.2 30.5
Noisy-AV (Noise) 2.5 83.0 49.3 24.9 12.7 15.2 14.2 15.6 31.9 27.7
Table 8: Language-wise WER of AVSR with different types of noise of Music and Noise from MUSAN dataset. The noise robustness of XLAVS-R is still the most effective with different kinds of noises.
@@ Appendix A Generalization to Different Types of Noise
Clean environment, Test WER (%) $\downarrow$
@@
Noisy environment, Test WER (%) $\downarrow$
@@
Clean environment, Test WER (%) $\downarrow$
@@
Noisy environment, Test WER (%) $\downarrow$
@@
Appendix B Future Works: Visual-only Experiment
The main focus of this research is on having the best noise-robust speech recognition and translation performance by leveraging stronger audio components and visual information, rather than visual-only input. Therefore, the presented AV models may not provide significant visual-only speech recognition capabilities, especially in non-English languages. The results of visual-only experiment in English are 52.4 WER for AV-HuBERT and 55.9 WER for u-HuBERT, and non-English results are around 90-100 WER. We plan to conduct comprehensive VSR experiment with XLAVS-R and further explore how to develop zero-shot visual speech recognition (VSR) ability while maintaining state-of-the-art audio-visual performances as a future work.